OpenStudy (anonymous):

How many solutions exist for the system, x + y = 1 and y = -x + 1? no solution one solution infinitely many solutions

OpenStudy (frostbite):

ok then we do it here? :P

OpenStudy (anonymous):

yeah sorry

OpenStudy (frostbite):

Okay... you do the work and I guide you: Rewrite: y = -x + 1

OpenStudy (anonymous):

what d you mean rewrite it its already in the form i need it right?

OpenStudy (frostbite):

Sure? what if I do this? \[x+y=1 \leftarrow \rightarrow y=1-x=-x+1\]

OpenStudy (frostbite):

The two equations are the same :)

OpenStudy (anonymous):

yes but what does that have to do with it?

OpenStudy (anonymous):

oooooooh *facepalm

OpenStudy (frostbite):

well a solution between two lines is when they interact... what if two graphs are on each other? ^^

OpenStudy (anonymous):

its no solutipon right?

OpenStudy (frostbite):

Nope there are infinitely many solutions.

OpenStudy (anonymous):

hu? oh so they would only be no soution if they were parallel?

OpenStudy (frostbite):

True ^^

OpenStudy (frostbite):

And I show you why now:

OpenStudy (anonymous):

k

OpenStudy (anonymous):

yeah i get oit thanks so much!